GLOBAL FIT ANALYSIS INCLUDING THE ν_{9} + ν_{4} − ν_{4} HOT BAND OF ETHANE: EVIDENCE OF IN-TERACTION WITH THE ν_{12} FUNDAMENTAL

J.R. COOPER, \underline{N. MOAZZEN-AHMADI}, Department of Physics and Astronomy, 2500 University Drive NW, Calgary, Alberta T2N 1N4, Canada.

The ν_{9} (methyl rock) fundamental band of ethane occurs in the 12 μm region. It is the strongest band of ethane in a terrestrial window and is commonly used for the identification of ethane in the Jovian planets. The ν_{9} + ν_{4} − ν_{4} band occurs in the same region; neither can be analysed as an isolated band, since both are embedded in the torsional bath of the ground vibrational state. We report here a global fit model including data from the ν_{3} (C-C stretch) fundamental, the ν_{4}, 2ν_{4} − ν_{4}, and 3ν_{4} torsional transitions, the ν_{9} fundamental, and its ν_{9} + ν_{4} − ν_{4} hot band. The combined data set of more than 4000 lines has been fitted to within experimental uncertainty using a model Hamiltonian which includes many terms which were important in analyzing the corresponding hot band of CD_{3}CD_{3}. We have therefore identified Coriolis interactions with the ground vibrational state (gs) for both the ν_{9} and ν_{9} + ν_{4} − ν_{4} bands, as well as gs/ν_{3} Fermi coupling, and have determined several barrier-dependence terms which were not previously accessible. More remarkably, we have also identified a new, resonant interaction with the ν_{12} fundamental which – unlike terms included in previous models – is off-diagonal in the torsional quantum number σ.